# Distilling heterogeneous treatment effects: Stable subgroup estimation in causal inference

### v1

@@ 3.3 Consistency of subgroup estimation: General setting with multiple rules
where $B_{r}=\{j_{r},\ldots,j_{r_{g}}\}$ denotes the set of remaining relevant subgroup feature indices, $A_{r}=\{1,\ldots,p\}\setminus B_{r}$ denotes the set of irrelevant subgroup feature indices, $s^{(a)}$ and $s^{(b)}$ correspond to the optimal split points ,conditional on $\mathcal{G}_{g}^{(r-1)}(X)=1$, for $X^{(a)}$ and $X^{(b)}$ respectively, and $\mathcal{G}^{(0)}_{g}(X)\equiv 1$.
Informally, Assumption (3.5) states that given the previously constructed rules, whether a rule is relevant or not must be distinguishable through the loss function. Moreover, $\delta$ represents the minimum difference at the population-level of the loss associated with splitting on a relevant and irrelevant covariate, conditional on the previous splits. This is intuitively similar to the irrepresentable condition for the LASSO (Zhao and Yu, 2006; Meinshausen and Bühlmann, 2006; Zou, 2006) or its variants (Javanmard and Montanari, 2013), which states that irrelevant features cannot be too correlated with relevant ones to achieve model selection consistency. If the losses between relevant and irrelevant rules are tied, then this means the tree will be unable to distinguish between what are relevant and irrelevant rules.
Assumption (3.5) ensures that the subgroups follow a problem with optimal substructure, such that a greedy search approach to constructing the second-stage decision tree learner is valid. In settings where Assumption (3.5) is not met, standard greedy tree algorithms such as CART can fail to recover the optimal partitioning even with infinite amounts of data (see Tan et al., 2024 for a formal discussion). Recent work has introduced alternative algorithms to construct optimal trees under weaker assumptions, but at a much higher computational cost (e.g., Hu et al., 2019; Lin et al., 2020), and are often restricted to specific settings such as binary features. Moreover, several studies have demonstrated the out-of-sample performance gains of optimal trees is only 1-2%, and the expected depth of greedy trees remains close to that of optimal trees (e.g., van der Linden et al., 2024). Given the popularity of greedy search approaches and the computational burden of many optimal tree construction approaches, we focus our discussion on settings in which greedy search algorithms can feasibly construct optimal subgroups.
@@
Let Assumptions (3.1)-(3.3) hold for all relevant covariates, conditional on the previously select decision rules. Additionally, assume bounded moments (Assumption (3.4)) and separability (Assumption (3.5)). Then, the difference between the estimated subgroup $\hat{\mathcal{G}}_{g}(X_{i})$ using CDT and the true subgroup $\mathcal{G}_{g}(X_{i})$ is upper bounded by the following:
@@
for all groups $g\in\{1,\ldots,G\}$, where $r_{g}$ corresponds to the number of rules in the $g^{th}$ subgroup, $k=\operatorname*{arg\,max}_{k^{\prime}}\left\{\left\lvert\hat{s}_{n}^{(k^{\prime})}-s^{(k^{\prime})}\right\rvert+\mathbb{E}_{n}\left[\ell(\hat{\tau}^{d}_{i};X^{(k^{\prime})},s^{(k^{\prime})})\right]-\mathbb{E}\left[\ell(\hat{\tau}^{d}_{i};X^{(k^{\prime})},s^{(k^{\prime})}\right]\right\}$, $M$ is a finite positive constant, and $\delta$ corresponds to the constant in the separability condition. Applying Proposition 3.1:
$\hat{\mathcal{G}}_{g}(X_{i})\stackrel{{\scriptstyle p}}{{\rightarrow}}\mathcal{G}_{g}(X_{i}).$
The results of Theorem 3.1 follow from first noting that the estimated splits will converge in probability to the optimal split (i.e., Lemma 3.1), and then showing that the within-sample loss will converge in probability to the true population-level loss. (See Appendix B for proof.) In particular, the rate of convergence of $\hat{\mathcal{G}}_{g}(X_{i})$ to $\mathcal{G}_{g}(X_{i})$ is dependent on (1) the degree of separability in the losses between relevant and irrelevant rules (i.e., $\delta$), and (2) the underlying convergence rate of the sample splits to the true optimal splits (i.e., the rate at which $\left\lvert\hat{s}_{n}^{(k)}-s^{(k)}\right\rvert\to 0$). An immediate implication of Theorem 3.1 is that not only will the tree consistently split at the correct cutoff points, but it will also correctly select the important features that should be used in the splitting criteria.
@@
Consider a simple data generating process, in which $X_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\text{MVN}(0,I)$ and $\tau_{i}=2\cdot\mathbf{1}\left\{X_{i}^{(1)}>0\right\}-\mathbf{1}\left\{X_{i}^{(2)}<-0.5\right\}+\epsilon_{i}$, where $\epsilon_{i}\sim N(0,\sigma_{\tau}^{2})$ (details in Section 5). For both CDT and causal trees, we calculate the total number of subgroups estimated, the number of true positive features (i.e., correctly estimated subgroup features), and number of false positive features (i.e., incorrectly estimated subgroup features). The number of estimated subgroups using causal tree increases linearly as a function of the sample size, while CDT correctly recovers the true underlying subgroups. Figure 3 visualizes the results.
Figure 3: Performance of causal trees and CDT (using causal forests as the teacher model), measured via the (A) number of estimated subgroups as well as (B) the number of true positive and (C) number of false positive features used in the estimated subgroups, under the additive subgroup data-generating process with linear covariates simulation (detailed in Section 5). The oracle number of subgroups, true positives, and false positives are shown as dashed gray lines. Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
To summarize, we have shown that under a set of regularity assumptions and a well-specified teacher model, CDT is able to consistently recover the set of optimal subgroups. Furthermore, CDT improves the efficiency of subgroup recovery over existing tree-based approaches by first de-noising the underlying data using the teacher model. The act of distilling thus allows researchers to leverage the interpretable outputs of decision trees in constructing subgroups, while offsetting the instabilities that usually adversarially impact tree estimation.
@@ 3.4 Estimating subgroup average treatment effects
In the following subsection, we consider how to estimate the subgroup average treatment effects. For simplicity, we will focus on an experimental setting, with extensions for estimating the subgroup average treatment effect in observational settings in Appendix A.4.
More formally, we assume that there is random treatment assignment in the study, such that the treatment indicator $Z$ is independent of $Y(1),Y(0)$, and $X$.
@@
Furthermore, we assume positivity, such that all units have a non-zero probability of receiving treatment.
@@
There exists some constant $0<\eta\leq 0.5$ such that:
$\eta<\Pr(Z=1\mid X)\leq 1-\eta.$
@@
We will take a finite-sample perspective, where we consider only the variation that arises from the treatment assignment process. Following existing literature, we condition on a set of valid randomizations—i.e., the set of random treatment allocations, such that there are treatment and control units in each subgroup $g\in\{1,\ldots,G\}$. [^2] Because of the sparsity induced by the causal distillation trees, we expect that the total number of subgroups $G$ should be small, relative to the sample size $n$. As such, we anticipate this restriction will have little impact on the results in practice. See Schochet (2024) and Miratrix et al. (2013) for more discussion. With some abuse of notation, we suppress the explicit conditioning in the results presented in the following section.
Because we are using sample splitting and honest estimation to estimate the subgroups, the estimated subgroups can be treated as a priori defined strata within the study. As such, under Theorem 3.1, the subgroup difference-in-means estimator, defined in Equation (6), provides a consistent estimate of the true subgroup average treatment effects.
Theorem 3.2 (Bias and Variance of Subgroup Difference-in-Means Estimator).
Under Assumptions (3.1)-(3.7), the bias of the difference-in-means across the units in the estimated subgroup $\hat{\mathcal{G}}_{g}(X_{i})$ (i.e., $\hat{\tau}^{(g)}$, as defined in Equation (6)) is upper bounded by the following:
$\lvert\mathbb{E}\left(\hat{\tau}^{(g)}\right)-\tau^{(g)}\rvert$ $\leq\frac{1}{n_{g}}\sum_{i=1}^{n}\left\lvert\hat{\mathcal{G}}_{g}(X_{i})-\mathcal{G}_{g}(X_{i})\right\rvert\Big{\lvert}Y_{i}(1)-Y_{i}(0)\Big{\rvert},$
where $n_{g}$ is the number of samples in subgroup $g$. Moreover, the variance of $\hat{\tau}^{(g)}$ is
$\text{var}(\hat{\tau}^{(g)})=$ $\frac{1}{n_{g}}\left\{\beta_{g}(1)\text{var}_{g}(Y_{i}(1))+\beta_{g}(0)\text{var}_{g}(Y_{i}(0))-\text{var}_{g}(\tau_{i})\right\}$
where $\beta_{g}(z)=\mathbb{E}\{1/\sum_{i=1}^{n}\hat{\mathcal{G}}(X_{i})\mathbbm{1}\{Z_{i}=z\}\}$ and $\text{var}_{g}\left\{Y(z)\right\}:=\mathbb{E}\left\{\frac{1}{n_{g}-1}\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})(Y_{i}(z)-\overline{Y_{g}(z)})^{2}\right\}$ for $z\in\{0,1\}$.
There are several takeaways to highlight from Theorem 3.2. First, we see that there is finite-sample bias that arises from slippage between the estimated subgroups and the true subgroups. The rate of convergence in $\hat{\tau}^{(g)}$ to $\tau^{(g)}$ will depend on how quickly $\hat{\mathcal{G}}_{g}(X_{i})$ converges to $\mathcal{G}_{g}(X_{i})$.
Second, from Theorem 3.2, we can construct a conservative variance estimator using the sample analogs for $\text{var}_{g}(Y_{i}(z))$ for $z\in\{0,1\}$:
@@
where $\hat{\beta}_{g}(z)=\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})\mathbbm{1}\{Z_{i}=z\}\right)^{-1}$ (i.e., the inverse of the proportion of units with treatment $Z_{i}=z$ in the $g$-th subgroup). Miratrix et al. (2013) show that the gap between $\hat{\beta}_{g}(z)$ and $\beta_{g}(z)$ can be upper bounded as a function of $1/n$. Asymptotically, this will be equivalent to the robust Huber–White variance estimator.
To consider the asymptotic distribution of the subgroup difference-in-means estimator, we must account for the fact that as the sample size increases, the underlying subgroup allocation will also change. We leverage recent work in developing central limit theorem results for finite populations (e.g., Schochet, 2024; Schochet et al., 2022; Li and Ding, 2017) to show that under mild regularity conditions, the subgroup difference-in-means estimator will be asymptotically normal.
@@
then the subgroup difference-in-means estimator will converge in distribution to a normal distribution:
@@
The results of Theorem 3.3 can be viewed as a special case of Schochet (2024), Theorem 1. The condition in (7) is a Lindeberg-type condition, which restricts the tail behavior of the underlying potential outcome distribution and is relatively weak. Importantly, both Theorem 3.2 and 3.3 highlight that the stability improvements we obtain from distillation affect not only our ability to better recover estimates of the subgroups, but also our ability to perform inference.
@@
Following Theorem 3.3, we can apply standard nonparametric tests of treatment effect heterogeneity. See Appendix A for details. In settings when researchers are interested in evaluating whether the largest treatment effect across the subgroups is statistically significantly different than the other treatment effects, we caution that a different test is needed due to post-selection bias. See Wei et al. (2024) for more discussion on valid re-sampling procedures for inference in these settings.
4 A stability-driven approach for teacher model selection
@@
In Section 3, we showed that the benefits of distillation arise from the teacher model’s ability to construct a meaningful basis representation of the individual-level treatment effect using the covariate data. In practice, there are different metalearners that researchers must choose from to estimate the individual-level treatment effect.
Although many model selection procedures exist for choosing the best prediction model (e.g., Stone, 1974; Akaike, 1974; Schwarz, 1978; Therneau et al., 1997; Breiman et al., 1984), they cannot be directly used to assess the goodness-of-fit of our teacher models, which predict individual-level treatment effects — quantities that are not observable in practice. Furthermore, even checking whether the predicted outcomes match the observed outcomes is insufficient to guarantee that we have modeled the treatment effects well (Künzel et al., 2019).
@@
In the following section, we develop a novel model selection procedure, designed specifically for subgroup estimation, that selects a teacher model based on its ability to stably reconstruct subgroups. Stable reconstruction refers to being able to estimate similar subgroups, given various perturbations in the underlying data generating process (i.e., sampling error). Following previous studies (e.g., Yu, 2013; Yu and Kumbier, 2020), we argue that teacher models that are able to construct similar subgroups across these different perturbations are likely more appropriate models for the data.
4.1 Quantifying Subgroup Similarity
@@
Before introducing the model selection procedure, we must first quantify what it means for two subgroups to be ‘similar’ and develop a novel measure, termed the Jaccard Subgroup Similarity Index (SSI). Informally, SSI considers two different sets of estimated subgroups $\{\hat{\mathcal{G}}^{(1)},\hat{\mathcal{G}}^{(2)}\}$, and computes the proportion of units that belong to the same subgroup across both partitions. If the units are similarly grouped in the two different partitions, then $\hat{\mathcal{G}}^{(1)}$ and $\hat{\mathcal{G}}^{(2)}$ are similar, and SSI is high. If the units are grouped differently, then $\hat{\mathcal{G}}^{(1)}$ and $\hat{\mathcal{G}}^{(2)}$ are less similar, and SSI will be low.
@@
To formalize this, let $\hat{\mathcal{G}}^{(1)}=\{\hat{\mathcal{G}}^{(1)}_{1}(X),\ldots,\hat{\mathcal{G}}^{(1)}_{G}(X)\}$ and $\hat{\mathcal{G}}^{(2)}=\{\hat{\mathcal{G}}^{(2)}_{1}(X),\ldots,\hat{\mathcal{G}}^{(2)}_{G}(X)\}$ denote two different sets of partitionings into subgroups. Then, for $k\in\{1,2\}$, define the matrix $C^{(k)}\in\mathbbm{R}^{n\times n}$ which denotes whether or not two distinct units belong in the same subgroup in $\hat{\mathcal{G}}^{(k)}$:
$C_{ij}^{(k)}=\begin{cases}1&\text{if }X_{i}\text{ and }X_{j}\text{ belong to the same subgroup in }\hat{\mathcal{G}}^{(k)}\text{ and }i\neq j,\\ 0&\text{otherwise,}\end{cases}$
for $i,j=1,\ldots,n$. We define the Jaccard Subgroup Similarity Index (denoted as SSI) as
@@
where $N_{qr}(\mathcal{H})=\lvert\{(i,j)\in[n]\times[n]:C^{(1)}_{ij}=q,\,C^{(2)}_{ij}=r,X_{i}\in\mathcal{H}(X)\}\rvert$ for each subgroup $\mathcal{H}\in\hat{\mathcal{G}}^{(1)}\cup\hat{\mathcal{G}}^{(2)}$ and $q,r=0,1$. Note that $N_{11}(\mathcal{H})$ is the number of sample pairs that belong to the same subgroup in both partitions ($\hat{\mathcal{G}}^{(1)}$ and $\hat{\mathcal{G}}^{(2)}$) but conditioned on the sample pair belonging to subgroup $\mathcal{H}(X)$, while $N_{01}(\mathcal{H})+N_{10}(\mathcal{H})$ is the number of sample pairs that belong to subgroup $\mathcal{H}(X)$ in one partition but two different subgroups in the other partition. The subgroup similarity index will be bounded between 0 and 1, such that when SSI is close to 1 (that is, when $N_{11}(\mathcal{H})$ is much larger than $N_{01}(\mathcal{H})+N_{10}(\mathcal{H})$), this implies the two sets of subgroups are similar. Figure A1 illustrates the computation of SSI for an example subgroup $\mathcal{H}\in\hat{\mathcal{G}}^{(1)}\cup\hat{\mathcal{G}}^{(2)}$.
@@
SSI is similar to the Jaccard index (Jaccard, 1901) used in comparing similarities across clusters (Ben-Hur et al., 2001). However, unlike the standard Jaccard index, which weights individual units equally, SSI re-weights units to give equal weight to each subgroup. This is necessary to penalize subgroups that comprise of very few units, which are typically highly unstable. [^3] We can consider the extreme setting, in which a subgroup contains only a single individual. An estimated subgroup with only one unit would have very little impact on the standard Jaccard index, even though in practice, such a subgroup would not be substantively meaningful and would be undesirable.
4.2 Teacher Model Selection Procedure
@@
Using SSI, researchers can then consider how similar the estimated subgroups are across repeated data samples. To mimic repeated sampling, we propose a bootstrapping procedure and detail the proposed teacher model selection procedure in Algorithm 1. At a high-level, for each candidate teacher model, researchers can use SSI to evaluate the stability of the estimated subgroups, obtained using the given candidate teacher model, across the different bootstrapped iterations. The teacher model that produces the highest SSI across the different bootstrapped iterations corresponds to the most stable teacher model and should be selected.
The teacher model selection procedure requires choosing a desired tree depth $d$, or equivalently, the number of rules in the estimated subgroups. As an example, a subgroup in a depth $d=2$ tree could be $\mathcal{G}_{g}(X)=\mathbbm{1}\{X_{\text{gender}}=\text{Female}\}\cdot\mathbbm{1}\{X_{\text{age}}<35\}$ while $\mathcal{G}_{g}(X)=\mathbbm{1}\{X_{\text{gender}}=\text{Female}\}\cdot\mathbbm{1}\{X_{\text{age}}<35\}\cdot\mathbbm{1}\{X_{\text{height}}>66\}$ is a subgroup in a depth $d=3$ tree. Domain expertise can sometimes dictate the choice of $d$. However, when researchers do not have a strong substantive prior for $d$, researchers can repeat the above procedure for multiple choices of $d$ and select the teacher model that yields the highest subgroup Jaccard stability scores across a majority of these $d$’s. We demonstrate the effectiveness of this approach for teacher model selection in Section 5.3.
1 for $l=1,\ldots,L$ do
2      Fit $\mathcal{M}^{l}$ on $\mathcal{D}^{train}$ to obtain estimates of the individual-level treatment effects $\hat{\tau}^{l}(X_{i})$ for each unit in $\mathcal{D}^{train}$ (via lines 2-16 in Algorithm 2)
3      for $b=1,\ldots,B$ do
4           for $t=1,2$ do
5                Get bootstrap sample of $\{(X_{i},\;\hat{\tau}^{l}(X_{i})):i\in\text{training unit}\}$
6                Fit decision tree $m$ on the bootstrapped $X_{i}$’s to predict the bootstrapped $\hat{\tau}^{l}(X_{i})$’s $\rightarrow$ estimated decision tree $\hat{m}^{(b_{t})}$
8
9                end for
11
12                end for
13
14                end for
Algorithm 1 Selecting the teacher model for CDT

### v2

@@ 3.3 General setting with multiple rules
where $B_{r}=\{j_{r},\ldots,j_{r_{g}}\}$ denotes the set of remaining relevant subgroup feature indices, $A_{r}=\{1,\ldots,p\}\setminus B_{r}$ is the set of irrelevant subgroup feature indices, $s^{(a)}$, $s^{(b)}$ are the optimal split points conditional on $\mathcal{G}_{g}^{(r-1)}(X)=1$ for $X^{(a)}$, $X^{(b)}$ respectively, and $\mathcal{G}^{(0)}_{g}(X)\equiv 1$.
Informally, Assumption (3.5) states that given the previously constructed rules, whether a rule is relevant or not must be distinguishable through the loss function. The constant $\delta$ represents the minimum difference of the population-level loss associated with splitting on a relevant and irrelevant covariate, conditional on the previous splits. This is intuitively similar to the irrepresentable condition for the Lasso (Zhao & Yu, 2006), which states that irrelevant features cannot be too correlated with relevant ones to achieve model selection consistency. If the losses between relevant and irrelevant rules are tied, then this means the tree will be unable to distinguish between what are relevant and irrelevant rules.
Assumption (3.5) ensures that the underlying subgroups have an optimal substructure so that a greedy second-stage decision tree learner is valid. If Assumption (3.5) is not met, standard greedy tree algorithms like CART can fail to recover the optimal partitions even with infinite amounts of data (see (Tan et al., 2024) for a formal discussion). Recent work has introduced alternative algorithms to construct optimal trees under weaker assumptions, but at a much higher computational cost, and are often restricted to specific settings such as binary features, with limited out-of-sample performance gains relative to greedy trees (e.g., Hu et al., 2019; van der Linden et al., 2024). Given the popularity of greedy search approaches and the computational burden of many optimal tree construction approaches, we focus our discussion on settings in which greedy search algorithms can feasibly construct optimal subgroups.
@@
Let Assumptions (3.1)-(3.3) hold for all relevant covariates, conditional on the previously select decision rules. Additionally, assume bounded moments (Assumption (3.4)) and separability (Assumption (3.5)). Then, for all groups $g\in\{1,\ldots,G\}$:
@@
where $k=\operatorname*{arg\,max}_{k^{\prime}}\left\{\left\lvert\hat{s}_{n}^{(k^{\prime})}-s^{(k^{\prime})}\right\rvert+\mathbb{E}_{n}\left[\ell(\hat{\tau}^{d}_{i};X^{(k^{\prime})},s^{(k^{\prime})})\right]-\mathbb{E}\left[\ell(\hat{\tau}^{d}_{i};X^{(k^{\prime})},s^{(k^{\prime})}\right]\right\}$, $M$ is a finite positive constant, and $\delta$ corresponds to the constant in the separability condition. Furthermore, applying Proposition 3.1, $\hat{\mathcal{G}}_{g}(X_{i})\stackrel{{\scriptstyle p}}{{\rightarrow}}\mathcal{G}_{g}(X_{i}).$
The results of Theorem 3.1 follow from first noting that the estimated splits will converge in probability to the optimal split (see Lemma 3.1), and then showing that the within-sample loss will converge in probability to the true population-level loss (see Appendix B) In particular, the rate of convergence of $\hat{\mathcal{G}}_{g}(X_{i})$ to $\mathcal{G}_{g}(X_{i})$ is dependent on (1) the degree of separability in the losses between relevant and irrelevant rules (i.e., $\delta$), and (2) the underlying convergence rate of the sample splits to the true optimal splits (i.e., the rate at which $\left\lvert\hat{s}_{n}^{(k)}-s^{(k)}\right\rvert\to 0$). An immediate implication of Theorem 3.1 is that not only will the tree consistently split at the correct cutoff points, but it will also correctly select the important features that should be used in the splitting criteria.
In practice, the total number of subgroups present is unknown, and can be treated as a tuning parameter. We show empirically that using standard pruning approaches result in consistent subgroup estimation with CDT (Appendix D). Furthermore, if researchers want to constrain the maximum number of rules in a subgroup, Theorem 3.1 guarantees that for a fixed number of rules $r$, CDT gives the optimal partitioning of that cardinality.
@@
Consider a simple data generating process, in which $X_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\text{MVN}(0,I)$ and $\tau_{i}=2\cdot\mathbf{1}\left\{X_{i}^{(1)}>0\right\}-\mathbf{1}\left\{X_{i}^{(2)}<-0.5\right\}+\epsilon_{i}$, where $\epsilon_{i}\sim N(0,\sigma_{\tau}^{2})$ (details in Section 5). We see that the number of estimated subgroups using causal tree increases linearly as a function of the sample size, while CDT correctly recovers the true underlying subgroups. Figure 2 visualizes the results.
Figure 2: Comparing the performance of causal trees versus CDT (using causal forests as the teacher model), measured via (A) number of estimated subgroups as well as (B) number of true positive and (C) number of false positive features used in the estimated subgroups. The oracle number of subgroups, true positives, and false positives are shown as dashed gray lines. Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
To summarize, we have shown that under a set of regularity assumptions and a well-specified teacher model, CDT is able to consistently recover the set of optimal subgroups. Furthermore, CDT improves the efficiency of subgroup recovery over existing tree-based approaches by first de-noising the underlying data using the teacher model. Distilling thus allows researchers to exploit the interpretable subgroup structure of decision trees, while offsetting the instabilities that usually adversarially impact tree estimation.
@@ 3.4 Estimating subgroup average treatment effects
Next, we consider how to estimate the subgroup ATEs. For simplicity, we focus on an experimental setting, with extensions for estimating the subgroup ATE in observational settings in Appendix A.3. More formally, we assume there is random treatment assignment in the study, such that the treatment indicator $Z$ is independent of $Y(1),Y(0)$, and $X$.
@@
We also assume positivity, such that all units have a non-zero probability of receiving treatment.
@@
There exists some constant $0<\eta\leq 0.5$ such that $\eta<\Pr(Z=1\mid X)\leq 1-\eta.$
@@
We will take a finite-sample perspective, where we consider only the variation that arises from the treatment assignment process. Following existing literature, we condition on a set of valid randomizations—i.e., the set of random treatment allocations, such that there are treatment and control units in each subgroup $g\in\{1,\ldots,G\}$. [^2] Because of the sparsity induced by the causal distillation trees, we expect that the total number of subgroups $G$ should be small, relative to the sample size $n$. As such, we anticipate this restriction will have little impact on the results in practice. See Schochet (2024) and Miratrix et al. (2013) for more discussion. With some abuse of notation, we suppress the explicit conditioning in the results presented next.
Because we are using sample splitting and honest estimation to estimate the subgroups, the estimated subgroups can be treated as a priori defined strata within the study. As such, under Theorem 3.1, the subgroup difference-in-means estimator, defined in Equation (4), provides a consistent estimate of the true subgroup ATEs.
Theorem 3.2 (Consistency and Variance of Subgroup Difference-in-Means Estimator).
Under Assumptions (3.1)-(3.7), the subgroup difference-in-means estimator will be an unbiased estimator for the sample subgroup ATE. Moreover, the variance of $\hat{\tau}^{(g)}$ is
$\text{var}(\hat{\tau}^{(g)})=\frac{1}{n_{g}}\left\{\beta_{g}(1)\text{var}_{g}(Y_{i}(1))+\beta_{g}(0)\text{var}_{g}(Y_{i}(0))-\text{var}_{g}(\tau_{i})\right\},$
where $\beta_{g}(z)=\mathbb{E}\{1/\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})\mathbbm{1}\{Z_{i}=z\}\}$ and $\text{var}_{g}\left\{Y(z)\right\}:=\mathbb{E}\left\{\frac{1}{n_{g}-1}\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})(Y_{i}(z)-\right.$ $\left.\overline{Y_{g}(z)})^{2}\right\}$ for $z\in\{0,1\}$.
If researchers are interested in the population-level subgroup ATE, then there will be some finite-sample bias that arises from slippage between the estimated subgroups $\hat{\mathcal{G}}(X)$ and the true optimal subgroups $\mathcal{G}(X)$; however, the subgroup difference-in-means estimator will still be a consistent estimator for the population-level subgroup ATE. (See Appendix A.4 for more discussion.)
From Theorem 3.2, we can also construct a conservative variance estimator using the sample analogs for $\text{var}_{g}(Y_{i}(z))$ for $z\in\{0,1\}$:
@@
where $\hat{\beta}_{g}(z)=\left(\frac{1}{n_{g}}\sum_{i=1}^{n}\hat{\mathcal{G}}_{g}(X_{i})\mathbbm{1}\{Z_{i}=z\}\right)^{-1}$. Miratrix et al. (2013) showed that the gap between $\hat{\beta}_{g}(z)$ and $\beta_{g}(z)$ can be upper bounded as a function of $1/n$. Asymptotically, this will be equivalent to the robust Huber–White variance estimator.
To consider the asymptotic distribution of the subgroup difference-in-means estimator, we must account for the fact that as the sample size increases, the underlying subgroup allocation will also change. We leverage recent work in developing central limit theorem results for finite populations (e.g., Schochet, 2024; Li & Ding, 2017) to show that under mild regularity conditions, the subgroup difference-in-means estimator will be asymptotically normal.
@@
then the subgroup difference-in-means estimator will converge in distribution to $N(0,1)$:
@@
The results of Theorem 3.3 can be viewed as a special case of Schochet (2024), Theorem 1. The condition in (5) is a Lindeberg-type condition, which restricts the tail behavior of the underlying potential outcome distribution and is relatively weak.
@@
Following Theorem 3.3, we can apply standard nonparametric tests of treatment effect heterogeneity (see Appendix A). If researchers are interested in evaluating whether the largest treatment effect across the subgroups is statistically significantly different than the other treatment effects, we caution that a different test is needed due to post-selection bias. See (Wei et al., 2024) for more discussion on valid re-sampling procedures for inference in these settings.
4 Stability-driven teacher model selection
@@
In Section 3, we showed that the benefits of distillation arise from the teacher model’s ability to construct a meaningful basis representation of the individual-level treatment effect using the covariate data. In practice, there are different metalearners that researchers must choose from to estimate the individual-level treatment effect. Traditional model selection procedures cannot be used to evaluate the goodness-of-fit of our teacher models, which predict individual-level treatment effects — quantities that are not observable in practice. Furthermore, checking whether the predicted outcomes match the observed outcomes is insufficient to guarantee that the treatment effects are modeled well (Künzel et al., 2019).
@@
We next introduce a novel model selection procedure, designed specifically for subgroup estimation. Our proposed procedure selects a teacher model based on its ability to stably reconstruct subgroups. Stable reconstruction refers to being able to estimate similar subgroups given various perturbations in the underlying data generating process. We consider more stable teacher models as being more appropriate models for the data, as they demonstrate greater robustness to sampling error (e.g., Yu, 2013; Yu & Kumbier, 2020).
@@
To begin, we must quantify what it means for two subgroups to be ‘similar’ and develop a novel measure, referred to as the Jaccard Subgroup Similarity Index (SSI). Informally, SSI considers two different sets of estimated subgroups $\{\hat{\mathcal{G}}^{(1)},\hat{\mathcal{G}}^{(2)}\}$ and computes the proportion of units that belong to the same subgroup across both partitions. If the units are similarly grouped in the two different partitions, then $\hat{\mathcal{G}}^{(1)}$ and $\hat{\mathcal{G}}^{(2)}$ are similar, and SSI is high. If the units are grouped differently, then $\hat{\mathcal{G}}^{(1)}$ and $\hat{\mathcal{G}}^{(2)}$ are less similar, and SSI will be low.
@@
To formalize, let $\hat{\mathcal{G}}^{(1)}=\{\hat{\mathcal{G}}^{(1)}_{1}(X),\ldots,\hat{\mathcal{G}}^{(1)}_{G}(X)\}$ and $\hat{\mathcal{G}}^{(2)}=\{\hat{\mathcal{G}}^{(2)}_{1}(X),\ldots,\hat{\mathcal{G}}^{(2)}_{G}(X)\}$ denote two different candidate sets of subgroups. Then, for $k\in\{1,2\}$, construct the adjacency matrix $C^{(k)}\in\mathbbm{R}^{n\times n}$ which denotes whether or not two distinct units belong in the same subgroup in $\hat{\mathcal{G}}^{(k)}$. Specifically, for $i,j=1,\ldots,n$, let $C_{ij}^{(k)}=1$ if $X_{i}$ and $X_{j}$ belong to the same subgroup in $\hat{\mathcal{G}}^{(k)}$ and $i\neq j$, and 0 otherwise. We then define the Jaccard Subgroup Similarity Index (SSI) as
@@
where $N_{qr}(\mathcal{H})=\lvert\{(i,j)\in[n]\times[n]:C^{(1)}_{ij}=q,\,C^{(2)}_{ij}=r,X_{i}\in\mathcal{H}(X)\}\rvert$ for each subgroup $\mathcal{H}\in\hat{\mathcal{G}}^{(1)}\cup\hat{\mathcal{G}}^{(2)}$ and $q,r=0,1$. $N_{11}(\mathcal{H})$ is the number of sample pairs that belong to the same subgroup in both partitions ($\hat{\mathcal{G}}^{(1)}$ and $\hat{\mathcal{G}}^{(2)}$), conditioned on the sample pair belonging to subgroup $\mathcal{H}(X)$. $N_{01}(\mathcal{H})+N_{10}(\mathcal{H})$ is the number of sample pairs that belong to subgroup $\mathcal{H}(X)$ in one partition but two different subgroups in the other partition.
@@
The subgroup similarity index is bounded between 0 and 1. When SSI is close to 1 (i.e., when $N_{11}(\mathcal{H})$ is much larger than $N_{01}(\mathcal{H})+N_{10}(\mathcal{H})$), this implies the two sets of subgroups are similar. Figure A2 illustrates the computation of SSI for an example subgroup $\mathcal{H}\in\hat{\mathcal{G}}^{(1)}\cup\hat{\mathcal{G}}^{(2)}$. SSI is similar to the Jaccard index (Jaccard, 1901) used to compare similarities across clusters (Ben-Hur et al., 2001). However, unlike the standard Jaccard index, which weights individual units equally, SSI re-weights units to give equal weight to each subgroup. This penalizes subgroups with very few units, which are typically highly unstable. [^3] We can consider the extreme setting, in which a subgroup contains only a single individual. An estimated subgroup with only one unit would have very little impact on the standard Jaccard index, even though in practice, such a subgroup would not be substantively meaningful and would be undesirable.
@@
Using SSI, we can then assess the similarity of estimated subgroups from each candidate teacher model across repeated data samples (e.g., using bootstrapping). The teacher model that produces the highest SSI across the different bootstrapped iterations corresponds to the most stable teacher model and should be selected (see Algorithm 2 for details).